You own a portfolio that has $1,800 invested in Stock A and $3,400 invested in Stock B. If the expected returns on these stocks are 11 percent and 18 percent, respectively, what is the expected return on the portfolio?(Do not round your intermediate calculations.)
Answer #1

porfeturn for =(w1 x r1) + (w2 x r2) + ... + (wn x rn) a PortfolioW1 = $1800/($1800 + $3400) = 0.3462

W2 = $3400/($1800 + $3400) = 0.6538

Expected return = 0.3462 * 11% + 0.6538 * 18%

Expected return = 15.58%